OpenStudy (anonymous):

Check my work please... \[\frac{ -10a ^{-4}b ^{-1} }{ 5a ^{-2}b ^{8} }\]

OpenStudy (anonymous):

-10 / 5 = -2 \[a ^{\left( -4 - (-2) \right)} = a ^{-2}\] \[b ^{\left( -1 - 8 \right)} = b ^{-9}\]

OpenStudy (anonymous):

Now how do I get the exponents to be positive in the final answer... here is the original question: Simplify the expression: \[\frac{ 10a ^{-4}b ^{-1} }{ 5a ^{-2}b ^{8} }\] Write your answer without negative exponents.

OpenStudy (anonymous):

@anyone

OpenStudy (phi):

\[ -2 a^{-2}b^{-9} \] to change the sign of the exponent, "flip" the base... if up top, put it in the bottom or vice versa

OpenStudy (phi):

so you get -2 over the other stuff

OpenStudy (anonymous):

So \[\frac{ -2 }{ a ^{2}b ^{9}}\]

OpenStudy (anonymous):

???

OpenStudy (phi):

yes, that is right.

OpenStudy (anonymous):

Awsome, thanks for your help!